$SPAD/src/input bern.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

f(t)==cos(t)/(1+sin(t)**2)
g(t)==sin(t)*cos(t)/(1+sin(t)**2)
c(t)==curve(f(t),g(t))
d:=draw(c(t)),t = -%pi..%pi, [title "Lemniscate of Bernoulli"])
close(d)

———-

References^{[1] nothing}

4